OpenStudy (dtan5457):

Solutions for 3cosx=cosx-1 domain=0<=x<2pi

OpenStudy (dtan5457):

should there be 4 solutions...?

OpenStudy (dtan5457):

1 per quadrant?

OpenStudy (dtan5457):

@dan815

OpenStudy (anonymous):

Nope! There should be 2 solutions tops

OpenStudy (dtan5457):

quadrant 2 and 3?

OpenStudy (anonymous):

Yep! On the right track.

OpenStudy (anonymous):

Do you need a hint?

OpenStudy (dtan5457):

ok

OpenStudy (anonymous):

3cosx=cosx-1 2cosx=-1 cosx = -1/2 right?

OpenStudy (dtan5457):

mhm

OpenStudy (anonymous):

So we are looking for the angle that gives us the point where x = -1/2

OpenStudy (anonymous):

Do you have a unit circle on you?

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

Look at the points where x is 1/2

OpenStudy (dtan5457):

2pi/3

OpenStudy (dtan5457):

is that theta 60?

OpenStudy (anonymous):

and there is another one

OpenStudy (dtan5457):

what angle is this on a 30-60-90?

OpenStudy (anonymous):

Dont get ahead of yourself! Find the other angle too!

OpenStudy (dtan5457):

i'm just asking what angle its on a 30-60-90..we won't have unit circles on a test.

OpenStudy (anonymous):

Yeah I memorized the circle. Um look at it like this

OpenStudy (dtan5457):

60, got it.

OpenStudy (dtan5457):

if the question was cosx=1/2?

OpenStudy (dtan5457):

4 solutions?

OpenStudy (anonymous):

We know this because the coordinates are -1/2 , sqrt(3)/2

OpenStudy (anonymous):

No 2 solutions

OpenStudy (anonymous):

Given that 0 > theta > 2pi

OpenStudy (anonymous):

Do you understand why?

OpenStudy (dtan5457):

has to be in the positive quadrants, right

OpenStudy (anonymous):

Yup